ROOT:5-x = x-3
do I first set both sides squared to get rid of the root? this again says check all proposed solutions which I don't understand.

Answer by edjones(7569)

sqrt(5-x)=x-3
5-x=x^2-6x+9 square each side.
x^2-5x+4=0 subtract the left side from the right.
(x-4)(x-1)=0 factor
x=4, x=1 These are the proposed solutions. Now you have to put them into the original equation and check to see if the solution is a true one or if it is extraneous.
.
sqrt(5-4)=4-3
sqrt(1)=1 true
.
sqrt(5-1)=1-3
2=-2 extraneous
.
Ed